A pine cone falls from the top of a 64-foot pine tree. How long will it take (in seconds) for the pine cone to reach the ground?
Use the equation h = -16t^2 + 64 where h is the height (in feet) above the ground t seconds after it falls from the tree.

Answer by Alan3354(69443)

A pine cone falls from the top of a 64-foot pine tree. How long will it take (in seconds) for the pine cone to reach the ground?
Use the equation h = -16t^2 + 64 where h is the height (in feet) above the ground t seconds after it falls from the tree.
--------------------------
h = -16t^2 + 64 = 0 (0 when it hits the ground)
16t^2 = 64
t^2 = 4
t = 2 seconds